### v1

# Filter-Banks for Ultra-Wideband Communications: Advantages and Design Challenges

Abstract
Recently, filter-bank multicarrier spread spectrum (FBMC-SS) has been proposed as a candidate waveform for ultra-wideband (UWB) communications. It has been noted that FBMC-SS is a perfect match to this application, leading to a trivial method of matching to the required spectral mask at different regions of the world. FBMC-SS also allows easy rejection of high-power interfering signals that may appear over different parts of the UWB spectral band. In this paper, we concentrate on the use of staggered multitone spread spectrum (SMT-SS) for UWB communications. SMT makes use of offset quadrature amplitude modulation (OQAM) to transmit data symbols over narrowband, overlapping subcarrier bands. This form of FBMC-SS is well-suited to UWB communications because it has good spectral efficiency and a flat power spectral density (PSD), resulting in good utilization of the UWB spectral mask. Additionally, we explore new methods for multi-coding that result in higher bit rates than previous FBMC-SS systems. Moreover, we study methods for equalizing the UWB multipath channel and cancelling narrowband interference. Excellent performance of the proposed methods are substantiated by presenting simulation results.
I Introduction
The idea of ultra-wideband (UWB) communications may be traced back to the spark gap radio transmission methods that were invented towards the end of the nineteenth century and the beginning of the twentieth century by Marconi and others. Spark gap radios, also known as impulse radios, were used for wireless telegraphy at the time. In the 1990s, when wireless technologies were booming and spectrum availability was becoming scarce, the possibility of using impulse radio UWB (IR-UWB) to develop spread spectrum radios that could coexist with existing (relatively, narrowband) radios with a minimum/acceptable interference to them was brought up, [1, 2]. The idea was well received by the Federal Communications Commission (FCC) which in 2002 allocated 7.5 GHz of spectrum for unlicensed use of UWB devices in the 3.1 to 10.6 GHz frequency band. This was to allow communications over short distances of up to 10 to 20 meters without licensing the spectrum.
The FCC had no waveform recommendation for UWB communications and instead specified that a UWB signal must occupy at least 500 MHz of bandwidth or have a bandwidth equivalent of at least 20 % of its center frequency, while satisfying a certain spectral mask for the transmit power spectral density (PSD) [3]. This spectral mask is defined by the regulatory bodies and varies in different regions of the world. Fig. 1 presents the spectral masks that have been defined for the USA (FCC), Europe (ECC), and Japan. As seen, these masks vary widely from place to place, hence, to be useable around the globe, a great degree of flexibility has to be included in any UWB system design.
Figure 1: UWB spectral mask in USA, Europe, and Japan.
After release of the 7.5 GHz of spectrum by the FCC, the two industrial organizations that emerged to establish standards for UWB, following IEEE802.15.3a, decided on waveforms that were not based on impulse radio. The first organization, UWB Forum, adopted a direct sequence (DS) spread spectrum-based UWB that was then referred to as DS-UWB, [4]. The second organization, the WiMedia Alliance, advocated for multiband orthogonal frequency division multiplexing (MB-OFDM) whose instantaneous bandwidth satisfied the minimum requirement of 500 MHz of the FCC definition for UWB, and a frequency hopping method was adopted to spread the transmit signal over a wider bandwidth, [5].
Despite significant efforts and investments by a large body of wireless industry (including, many established and hundreds of startup companies), by the beginning of 2006, the UWB technology development under IEEE802.15.3a had effectively stalled due to standardization issues. Reference [6] presents an excellent history of the first few years after the FCC allocation of the 7.5 GHz of spectrum for UWB communications and what happened at the end.
The subsequent follow up by the industry led to the IEEE802.15.4 series of standards, [7, 8]. The data rates in IEEE802.15.4 standards are significantly lower than the original rates that were envisioned in IEEE802.15.3a (up to 480 Mb/s). As opposed to IEEE802.15.3, IEEE802.15.4 standards adopted IR-UWB.
In a recent article, [3], filter-bank multicarrier (FBMC) has been introduced as a new candidate waveform for UWB communications. It has been noted that FBMC is a perfect match to the needs of UWB as it offers the following advantages:
  • Spectral mask adoption in different regions of the world is simplified through activation and deactivation of subcarriers.
  • The filter-bank structure lends itself to implementations that are tailored for parallel processing and avoid power-hungry high-speed analog-to-digital (A/D) and digital-to-analog (D/A) convertors.
  • The filter-bank structure allows adoption of a special receiver structure that blindly removes any interference. No other waveform in the broad area of digital communications offers this feature.
  • FB-UWB design supporting simultaneous operation of piconets is straightforward.
  • Ranging capability of FB-UWB is as straightforward as that of the IR-UWB.
The article [3] also provides detailed arguments supporting the above advantages of FBMC. In addition, [3] presents some numerical results showing FBMC based UWB can support transmission rates comparable to the best envisioned rates in the past literature. These numerical results are for the cases where either filtered multitone (FMT) or staggered multi-tone (SMT), [9], are used to implement an FBMC-SS waveform. However, details of any transceiver design has been left for future studies.
In this paper, we focus on a transceiver design based on SMT. SMT, in contrast to FMT, has a flat PSD and more efficiently utilizes the signal bandwidth. This is because the FMT subcarrier bands, by design, are non-overlapping. This leaves a null band between each pair of adjacent subcarrier bands. The presence of these spectral nulls, along with the need of satisfying the constraint imposed by the spectral mask, reduces the maximum power of the transmit signal, hence, leading to a reduced transmission range. On the other hand, by making use of offset quadrature amplitude modulation (OQAM), SMT allows overlapping the adjacent subcarrier bands, resulting in a flat PSD and hence better utilization of the UWB spectral mask [9]. Additionally, because the subcarrier bands have been overlapped, the spectral efficiency of SMT-SS is greater, thus allowing higher peak data rates and a full utilization of the available degrees of freedom.
This paper explores several important aspects of a UWB transmitter design when SMT is used for signal modulation. Using the IEEE802.15.4a UWB channel model [10], we show the typical received signal PSD levels for offices and industrial environments. This study provides some insight to the maximum data rates that may be available in the considered environments. We also explore several design opportunities for achieving a maximum data rate while facilitating practical methods of channel equalization and combating of high-power interferences that come from incumbent radios. Such interferences may be $30$ dB or more above the UWB spectrum.
The remainder of this paper is organized as follows. In Section II, we present a review of the most significant challenges in UWB system design and the previously proposed waveforms. In Section III, we present the FBMC-SS system model proposed in this paper and show a multi-coding method well-suited to the variations in received SNR typical of a UWB transmission. In Section IV, we present a design based on fast convolution and show how multi-tap equalization per subcarrier band can be implemented using a minimum mean square error (MMSE) based equalization scheme. In Section V, we demonstrate the performance of this UWB system design over simulated UWB multi-path channels with harsh interference. In Section VI we conclude.
II Challenges in UWB wireless systems
Though UWB has great potential for high-rate wireless communications, its adoption by the industry has remained limited. Early thoughts in the development of UWB systems envisioned data rates as high as $500$ Mbps. It was soon realized that these rates required sophisticated signal processing that was difficult to implement at the time. We believe this was one of the main reasons for the failure of 802.15.3 activities and the directing of the activities to the much lower data rates in 802.15.4 standards. In addition, the waveforms that were proposed by UWB Forum and WiMedia Alliance did not have sufficient flexibility to accommodate the much-needed signal processing steps. As discussed in [3] and further shown in this paper, filter-bank UWB is an enabler technology for high-rate UWB communications.
To look into challenges that any UWB system may face, we present some measured results which we have gathered using an actual radio. Following the parallel structure discussed in [3], we have constructed a UWB transceiver system with a bandwidth of $1.28$ GHz, using a set of Ettus X310 universal software radio peripherals (USRPs). This construction requires four X310 radios for the transmitter and the same number of radios for the receiver. Each USRP contains two independent RF channels that can be tuned to different carrier frequencies. Each channel is sampled at 200 MHz, with 160 MHz of useable bandwidth. Hence, the total bandwidth of $1.28$ GHz is obtained through a set of eight parallel radio chains (using four USRPs), including the respective A/D and D/A convertors. The USRPs can be synchronized by making use of an external master clock; one for each of the transmitter and receiver sides.
To get some insight to the challenges that one may face when implementing a UWB system, Fig. 2 presents a diagram showing the PSD of our equipment noise, along with the ranges of the received signal PSD for a set of line-of-sight (LOS) and non-line-of-sight (NLOS) channels based on the IEEE802.15.4a channel model, in an office setting and an industrial site. The equipment noise is obtained from the Ettus data sheets and also confirmed through our measurements, where the antenna port was terminated. The important point to note here is that the received signal is mostly at some negative SNR values. For NLOS channels, in particular, the SNR range is at $25$ dB or more below the equipment noise. At such a low SNR, even detection of the presence of a data packet may not be a straightforward task. This explains why in 802.15.4 standard document the synchronization part of each packet can be as long as $4096$ pilot symbols.
Figure 2: PSD of the equipment noise and the ranges of the received signal PSD as compared the internal noise generated by the radio equipment.
Another serious problem that any UWB system may encounter in practice is the presence of interference from incumbent radios. Fig. 3 presents a spectrogram of radio activities near an LTE base station tower over a period of 300 msec. Here, the equipment noise power is around $-90$ dBm. The UWB signal is at some lower level, hence, not observable here. Interference from incumbent radios is typically at a level $20$ dB or more above the equipment noise and is spread broadly across the time-frequency space. As seen, although there exist plenty of unused time-frequency points that a UWB system can benefit from, unless an effective receiver is used, the transmission rate may be greatly compromised because of interference from the incumbent radios.
Figure 3: A spectrogram of signal activities near an LTE tower over a bandwidth of $1.28$ GHz.
The above study of noise and signal PSD levels clearly shows that any effective implementation of a UWB system should make use of some form of a spread spectrum system. Moreover, it should be equipped with an effective mechanism for rejection of strong narrowband interferences from incumbent radios and other spectral activities in the environment. As noted in [3], filter-bank UWB is well-suited to addressing both of these challenges. In light of this, we present a brief review of FBMC modulation techniques and then show their use for implementation of a spread spectrum system that is matched to the needs of UWB communications. We also present simulation results to confirm the effectiveness of the proposed system in a UWB setup.
III Filter-Bank Multicarrier Spread Spectrum Modulation
As noted in Section I, SMT should be used for FBMC modulation in UWB systems. SMT allows for maximization of the transmit power (hence, maximizing the transmission range and/or data rate), while satisfying the spectral mask imposed by the FCC and other regulatory bodies around the world. SMT makes use of OQAM where the real and imaginary parts of each QAM data symbol are treated as two real-valued symbols and transmitted at the spacing of $T/2$, where $T$ is the QAM symbol interval. In addition, a $90^{\circ}$ phase-shift is applied to adjacent (real-valued) data symbols along time and frequency axes. This is to facilitate data recovery at the receiver, making use of what is referred to as real orthogonality in the basis set used in signal synthesis at the transmitter, [9].
To implement an SMT-based spread spectrum (SMT-SS) each QAM symbol is repeated across a number of $K$ subcarrier bands. This leads to a processing gain $G=10\log_{10}K$ dB. The choice of $K$ depends on the SNR at the receiver input and the desired SNR after despreading. For instance, if there is a LOS channel with an SNR of $-10$ dB at the receiver input and an SNR of $10$ dB is required after despreading, $K$ has to be set equal to $100$. This may be rounded up to $128$ for an effective multi-coding of $7$ bits for each of real and imaginary parts of each QAM symbols, assuming quadrature phase-shift keying (QPSK) signaling. This, for a transmission bandwidth of $1.28$ GHz, leads to a data rate $1402\times 7\times\frac{1,280}{128}=$ 140 Mbps. This data rate will, clearly, vary proportionately with the size of the transmission bandwidth. An example of one such multi-coding scheme in the context of FBMC is discussed in [11].
As a second example, consider the case of an NLOS channel with an SNR of $-30$ dB at the receiver input, i.e., $20$ dB lower than the case of LOS channel, above. To maintain the same output SNR of $10$ dB, this demands a processing gain which is $100$ times larger, i.e., $10,000$. Rounding this up to the nearest power of two, we get $K=2^{14}$ and a data rate $2\times 14\times\frac{1,280}{2^{14}}\approx 2.2$ Mbps.
Comparing the above two examples, one may notice that while the subcarrier bandwidth in the first example is $10$ MHz, in the second example, it reduces to $10/128=0.078125$ MHz. Accordingly, one may think of the need for a large variation of subcarrier bandwidth, to be able to cater to different needs of the UWB radios which may vary depending on the channel condition at any given time. This may be greatly simplified, if one chooses to keep the subcarrier bandwidth fixed to a minimum value and transmit through a set of parallel channels, to adjust the required processing gain and, accordingly, the data rate as explained below.
Consider a case where a total bandwidth of $W$ is divided into $M$ subbands, each with the bandwidth of $B=W/M$. If a given data symbol is spread across all the $M$ subbands, one will end-up with a symbol rate $B$ symbols per second and a processing gain $10\log_{10}M$ dB. Alternatively, if a data symbol sequence is partitioned into $K$ parallel subsequences and each subsequence is spread across $M/K$ subbands, there will be a total symbol rate $KB$ symbols per second and a processing gain $10\log_{10}\frac{M}{K}$ dB. In this way, one can maximize the transmission rate while achieving the required decoding SNR after despreading.
IV Fast Convolution Filter-Banks:
An Effective Tool for Implementation of UWB Transceiver Systems
Fast convolution (FC) is a classical method for efficient implementation of long finite impulse response (FIR) filters. More recently, the use of FC for implementation of filter-banks has been widely studied and many publications addressing different aspects of such systems have been published, e.g., [12, 13, 14]. For UWB, fast convolution architectures offer opportunities for low complexity, multi-tap channel equalization [15]. In this section, we explore a fast convolution based receiver design and show how a high-performance, low complexity channel equalization scheme can be developed that takes advantage of the diversity introduced by spreading data symbols over a number of subcarriers.
Fig. 4 shows a diagram of our proposed SMT fast convolution receiver. We design our fast convolution receiver based on the conventional overlap save fast convolution algorithm. First, the input sample stream is partitioned into overlapping blocks of length $N$, where the overlap between adjacent blocks is chosen to be $N_{o}$. The amount of overlap is chosen so that the portions of the output sample stream impacted by the effects of frequency domain filtering with the prototype filter and the multi-tap equalizer can be removed. The $N$-point DFT is taken of each block, yielding the frequency domain samples $\tilde{x}_{k}$, for $k=0,\ldots,N-1$. Note that since the adjacent subcarrier bands overlap, the samples at the DFT output share between the adjacent bands. For instance, in the left most part of Fig. 4, while the first and second blocks of samples are used for data recovery in the first subcarrier band, the second and third blocks of samples are used for data recovery in the second subcarrier band. Moreover, as discussed next, an equalization method that combines signals from different subcarrier bands while taking into account partial band interferers is possible.
Figure 4: Fast convolution structure for an SMT receiver.
Fractionally spaced, multi-tap equalizers for SMT systems have been discussed in many publications, with [16] providing a good introduction and summary. These approaches include linear and widely linear MMSE equalization techniques, embedded equalizers in FC-based receivers [15], and multi-stage interference cancellation approaches. In connection with our fast convolution architecture, we choose a frequency domain equalization and interference cancellation scheme that takes advantage of the spread spectrum nature of the proposed SMT-SS design.
Frequency domain equalizers have been discussed in a number of publications, [17, 15, 16]. They have been shown to offer low complexity per-tone equalization, allowing equalizer weights at different DFT tones to be calculated independently [17]. Additionally, here, we show taking partial band interference into account in a frequency domain equalizer is also possible and leads to a significant performance improvement.
In designing our channel equalizer and interference suppression scheme, we extend the MMSE designs of [15, 17]. Let the $n^{\text{th}}$ frequency domain sample in the $k^{\text{th}}$ subcarrier band corresponding to the $m^{\text{th}}$ data stream be denoted $\tilde{x}^{(m)}_{k,n}$. Also, let $c_{k,n}$ and $h_{k,n}$ denote the corresponding channel and matched filter gains respectively and let $\sigma^{2}_{k}$ denote the noise plus interference power over the $k^{\text{th}}$ subcarrier band. To arrive at an MMSE equalizer, we first apply a noise whitening step
$\tilde{x}^{(m)^{\prime}}_{k,n}=\frac{\tilde{x}^{(m)}_{k,n}}{\sqrt{\sigma_{k}^{2}}}.$ (1)
Once the noise power has been whitened, the multichannel, frequency domain equalizer of [17] can be directly applied. This leads to the respective equalizer tap weight
$\tilde{w}^{(m)^{\prime}}_{k,n}=\frac{\frac{h_{k,n}c^{*}_{k,n}}{\sigma_{k}}}{\sum_{l\in\mathcal{S}_{m}}\left(\left|\frac{h_{l,n}c_{l,n}}{\sigma_{l}}\right|^{2}+\left|\frac{h_{l,n^{\prime}}c_{l,n^{\prime}}}{\sigma_{l}}\right|^{2}\right)+\frac{1}{\text{SNR}^{^{\prime}}_{m}}},$ (2)
where $\mathcal{S}_{m}$ is the set of all subcarrier indices corresponding to the $m^{\text{th}}$ data stream, $n^{\prime}$ refers to the tone indices that alias to $n$, and $\text{SNR}^{^{\prime}}_{m}$ is the averaged SNR across the associated subcarrier bands, given by
$\text{SNR}^{^{\prime}}_{m}=\frac{1}{L}\sum_{l\in\mathcal{S}_{m}}\text{SNR}_{l},$ (3)
where $L$ is the number of subcarrier bands in a data stream.
Making use of (1) and (2), the equalized signal corresponding to the $m^{\text{th}}$ data stream at tone $n$ is given by
$\tilde{y}_{m,n}=\sum_{k\in\mathcal{S}_{m}}\tilde{w}^{(m)^{\prime}}_{k,n}\tilde{x}^{(m)^{\prime}}_{k,n}.$ (4)
Alternatively, this may be rewritten as
$\tilde{y}_{m,n}=\sum_{k\in\mathcal{S}_{m}}\tilde{w}^{(m)}_{k,n}\tilde{x}^{(m)}_{k,n},$ (5)
where $\tilde{x}^{(m)^{\prime}}_{k,n}$ is replaced by $\tilde{x}^{(m)}_{k,n}$ and, accordingly, we use the equalizer tap weight
$\tilde{w}^{(m)}_{k,n}=\frac{\frac{h_{k,n}c^{*}_{k,n}}{\sigma_{k}^{2}}}{\sum_{l\in\mathcal{S}_{m}}\left(\left|\frac{h_{l,n}c_{l,n}}{\sigma_{l}}\right|^{2}+\left|\frac{h_{l,n^{\prime}}c_{l,n^{\prime}}}{\sigma_{l}}\right|^{2}\right)+\frac{1}{\text{SNR}^{^{\prime}}_{m}}},$ (6)
This, effectively, is a maximum ratio combining (MRC) equalizer. It gives lower weights to the bands that are highly interfered and/or are affected by a lower channel gain.
V Simulations
TABLE I: Simulation Configurations
[…]
Channel
Type
Bit Rate
BW (MHz)
Subcarrier
BW (KHz)
LDPC Code
Rate
LOS 456.25 5120 311.9 5/7
NLOS 0.08 500 243.6 1/3
We demonstrate the performance of the proposed fast convolution SMT receiver design through a number of simulations. In particular, we compare the performance of a system that uses the method discussed in Section IV, and a method that performs MMSE equalization and combining, but without considering the received noise PSD, i.e., when the noise whitening step is removed. This method is representative of a system design that uses a per-tone channel equalizer but does not consider narrowband interference suppression.
In our simulations, we show two parameterizations of the design, one that is well-suited to low rate transmissions in NLOS channels and the other that is well-suited to high rate transmissions in LOS channels. Key parameters of the two configurations are given in Table I. These parameters are chosen based on the anticipated received signal SNR for typical LOS and NLOS scenarios over a range of 33338888 meters. Received signal strength is found using the IEEE802.15.4 channel model and the noise power is chosen based on the noise figure of our development hardware, where it has been found that the SNR in typical LOS scenarios over this range is between $-10$ dB and $+3$ dB. For the case of NLOS channels, the typical received SNR is between $-42$ dB and $-23$ dB.
All simulations use full receiver channel state information (CSI) and are performed in a simulated primary user interference environment. The interference environment is characterized based on an extensive study of the IEEE802.15.4 operating bands. It is found that typical interference events range from 555540404040 dB above the receiver noise level. For each simulated data block, a number of interfering signals each with a bandwidth of 20 MHz are generated with a random carrier frequency and power spectral density level between $5$ and $40$ dB above the noise power level and added to the received signal. As the system bandwidth grows, the number of potential interfering signals also increases. For this reason, we chose 10 interfering signals for the high rate system, occupying a bandwidth of $5120$ MHz, and 4 interferers for the low rate system, occupying a bandwidth of $500$ MHz.
Bit-error-rate (BER) curves for low rate systems operating over the NLOS office channel are shown in Fig. 5. As can be seen, the bit error rate falls off around an SNR of $-36$ dB when narrowband interference suppression is applied. Lower SNR performance can be achieved through either a bandwidth expansion or a lower transmission rate. BER curves for the high rate system are shown in Fig. 6. It can be seen that by suppressing partial band interference, a rate of $456.25$ Mbps can be achieved for typical SNR values seen in LOS channels. On the other hand, when no interference suppression is applied, the BER remains very high. Also, it should be noted that the data rates in the order of $500$ Mbps require a bandwidth that exceeds an order of magnitude larger than the typical bandwidth of $500$ MHz, in use in current UWB systems.
Figure 5: Low rate curves for the NLOS industrial and office channels.
Figure 6: High rate curves for the LOS industrial and office channels.
VI Conclusion
An exploration of filter-banks for communications in realistic UWB operating environments was presented. First, through a hardware setup, we studied the interference environments that UWB systems are likely to encounter. The conclusion drawn from this study was that UWB systems are likely to encounter a large number of narrowband interferers at random locations in the band of transmission, at a level $5$ to $40$ dB above the background noise. It was then argued that an effective UWB receiver should be equipped with a flexible mechanism for filtering such interferers and, following [3], it was noted that a transceiver implementation based on filter-banks may provide the best solution. Moreover, by making use of the IEEE802.15.4 channel model, we found that, in most cases, a received UWB signal falls at a level below the noise floor, sometimes as low as $-42$ dB below the noise floor. We thus proposed and studied the use of a filter-bank spread spectrum system for UWB communications. The effectiveness of the proposed system was studied through computer simulations over both LOS and NLOS UWB channels.
The results demonstrated in this paper show that SMT, combined with receiver architecture shown in Section IV is a highly effective UWB communication method. The effectiveness of this method was demonstrated in the presence of harsh interference, over standardized LOS and NLOS UWB channel models with typical SNRs for both cases. Additionally, due to the spectral partitioning properties of filter-banks and the parallel architecture proposed in [3], these algorithms can be implemented with moderate sampling rates.
Acknowledgements
This research made use of Idaho National Laboratory’s High Performance Computing systems located at the Collaborative Computing Center and supported by the Office of Nuclear Energy of the U.S. Department of Energy and the Nuclear Science User Facilities under Contract No. DE-AC07-05ID14517.
This manuscript has in part been authored by Battelle Energy Alliance, LLC under Contract No. DE-AC07-05ID14517 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the paper for publication, acknowledges that the United States Government retains a nonexclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. STI Number: INL/JOU-24-77579.